### v2

# M2NO: Multiresolution Operator Learning with Multiwavelet-based Algebraic Multigrid Method

Abstract
Solving partial differential equations (PDEs) effectively necessitates a multi-scale approach, particularly critical in high-dimensional scenarios characterized by increasing grid points or resolution. Traditional methods often fail to capture the detailed features necessary for accurate modeling, presenting a significant challenge in scientific computing. In response, we introduce the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), a novel deep learning framework that synergistically combines multiwavelet transformations and algebraic multigrid (AMG) techniques. By exploiting the inherent similarities between these two approaches, M2NO overcomes their individual limitations and enhances precision and flexibility across various PDE benchmarks. Employing Multiresolution Analysis (MRA) with high-pass and low-pass filters, the model executes hierarchical decomposition to accurately delineate both global trends and localized details within PDE solutions, supporting adaptive data representation at multiple scales. M2NO also automates node selection and adeptly manages complex boundary conditions through its multiwavelet-based operators. Extensive evaluations on a diverse array of PDE datasets with different boundary conditions confirm M2NO’s superior performance. Furthermore, M2NO excels in handling high-resolution and super-resolution tasks, consistently outperforming competing models and demonstrating robust adaptability in complex computational scenarios.
1 Introduction
Partial differential equations (PDEs) are fundamental to modeling a wide range of phenomena across science and engineering [13], including fluid dynamics [28, 24], material stress analysis [32], and atmospheric conditions. Traditionally, solving these equations requires specific numerical methods, designed to address individual problem characteristics, often resulting in significant computational efforts [17]. Recent developments in deep learning have introduced alternative approaches to solving PDEs [25, 22, 23, 8, 21], featuring methods like Fourier Neural Operators (FNO) [22], Galerkin Transformers (GT) [8] and Deep Operator Networks (DeepONet)[23]. These techniques leverage neural network architectures to approximate the mappings between parameters and solutions in infinite-dimensional spaces, efficiently addressing both forward and inverse problems in PDEs. Among these techniques, the Fourier Neural Operators (FNO) [22, 20] stand out as a particularly significant development. FNO integrates Fast Fourier Transforms within a neural network framework, effectively learning direct mappings between infinite-dimensional function spaces. This capability marks FNO as a powerful tool, offering a scalable and efficient alternative to traditional numerical methods for solving PDEs.
A primary limitation of the Fourier Neural Operator (FNO) is its reliance on basis functions generated by Fast Fourier Transforms (FFTs). These functions are localized in frequency but lack spatial resolution, leading to difficulties in managing complex boundary conditions and capturing local features in high-dimensional PDEs [2, 9]. In contrast, wavelets provide both spatial and frequency localization [9, 2, 3], which enhances the management of discontinuities and improves image pattern recognition over FFTs. Despite these advantages, the Multiwavelet Transformation (MWT) model, which utilizes wavelet kernels computed by four independent neural networks, faces its own set of challenges. While leveraging Multi-Resolution Analysis (MRA) to provide a detailed representation of both frequency and spatial information, MWT requires substantial computational resources during training and struggles with grid sizes that are not powers of two, which limits its generalizability and practical application [12].
This limitation in MWT serves as a segue into discussing the Algebraic Multigrid (AMG) method. Although AMG is highly effective in improving computational efficiency in numerical computing, it is traditionally constrained by its compatibility with only simple periodic boundary conditions. Moreover, AMG requires the pre-determination of mapping operators between grid points at each level, which poses further limitations on its generalizability and practicality [6, 35, 29, 4].
In response to these limitations, we introduce an innovative neural network architecture, the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), which synergistically combines the strengths of algebraic multigrid methods and multiresolution analysis [5, 11, 7]. Our contributions are designed to overcome the specific challenges faced by previous models:
  • Enhanced Operator Learning: M2NO merges algebraic multigrid and multiresolution techniques to facilitate adaptive learning across multiple resolutions. This integration not only enables automatic node selection but also significantly improves the handling of complex PDE coefficients and intricate boundary conditions.
  • Advanced Operator Design: We develop advanced prolongation and restriction operators based on multiwavelet theory. These operators are critical components of our neural operator framework, enabling efficient hierarchical decomposition that captures both global and local structural details within PDE solutions.
  • Empirical Validation: M2NO is rigorously tested across a variety of 1D and 2D PDE datasets under different boundary conditions. Our experiments confirm its superior accuracy and scalability, consistently outperforming traditional models at various resolution scales.
These innovations underscore our commitment to advancing PDE computational frameworks, offering robust solutions for complex mathematical challenges in scientific computing.
2 Algebraic Multigrid Method and Multiwavelet
In this section, we first provide a concise overview of the algebraic multigrid method [35, 6] and multiresolution analysis using multiwavelets [1, 9]. We then summarize the similarities between these approaches and illustrate how they inform our development of multiwavelet-based prolongation and restriction operators.
2.1 Algebraic Multigrid Method
The algebraic multigrid method is a well-grounded numerical technique for solving systems of algebraic equations that approximate partial differential equations (PDEs). Specifically, we address the problem of solving the linear equation system:
$A^{h}u^{h}=b^{h},$ (1)
where $A^{h}$ and $b^{h}$ are derived from the discretization of a differential equation on a grid $\Omega^{h}$, with $h$ representing the step size.
Given a prolongation operator, $I_{2h}^{h}$, where the superscript denotes the fine grid and the subscript denotes the coarse grid, alongside a restriction operator, $I_{h}^{2h}$, we can recursively define a multigrid method. In the two-level V-cycle method, we first perform a few relaxation steps (typically one or two) on the fine grid $\Omega^{h}$ to obtain an initial estimate $u^{h}$. Subsequently, we calculate the residual $r^{h}=b^{h}-A^{h}u^{h}$; restrict this residual to the coarse grid $\Omega^{2h}$, yielding $\Omega^{2h}$, $r^{2h}=I_{h}^{2h}r^{h}$; and solve the error equation
$A^{2h}e^{2h}=r^{2h}$ (2)
on the coarse grid. After solving, update the solution on the fine grid by setting
$u^{h}=u^{h}+I_{2h}^{h}e^{2h}$ (3)
and perform another few relaxation steps (typically one or two) on the fine grid. This procedure outlines the foundational steps of the V-cycle multigrid scheme, which is recursively defined, as shown in Algorithm 1. For a comprehensive discussion and theoretical grounding of the Algebraic Multigrid (AMG) methods, please refer to the Appendix A.
2.2 Multiwavelets
Wavelets are powerful tools for decomposing data, functions, or operators into different frequency components and analyzing them through scaling. They can form a complete orthonormal basis for $L^{2}(\mathbb{R})$. Due to their scaling properties, wavelet functions have time or space widths inversely related to their frequencies: they are narrow at high frequencies and broad at low frequencies. This characteristic ensures excellent localization in both the frequency domain and physical space, making wavelets particularly effective for analyzing phenomena across multiple scales.
A multiresolution analysis (MRA) using multiwavelets involves a sequence of closed subspaces, defined as:
$\mathbf{V}_{0}^{k}\subset\mathbf{V}_{1}^{k}\subset\ldots\subset\mathbf{V}_{n}^{k}\subset\ldots$
For $k=1,2,\ldots$, and $n=0,1,2,\ldots$, we define $\mathbf{V}_{n}^{k}$ as a space of piecewise polynomial functions with dimension $2^{n}k$, that satisfy certain conditions. The multiwavelet subspace $\mathbf{W}_{n}^{k}$, defined as the orthogonal complement of $\mathbf{V}_{n}^{k}$ in $\mathbf{V}_{n+1}^{k}$, satisfies:
$\mathbf{V}_{n}^{k}\oplus\mathbf{W}_{n}^{k}=\mathbf{V}_{n+1}^{k},\quad\mathbf{W}_{n}^{k}\bot\mathbf{V}_{n}^{k},$ (4)
and the dimension of $\mathbf{W}_{n}^{k}$ is of dimension $2^{n}k$. Therefore, the composition of spaces is:
$\mathbf{V}_{n}^{k}=\mathbf{V}_{0}^{k}\oplus\mathbf{W}_{0}^{k}\oplus\mathbf{W}_{1}^{k}\oplus\ldots\oplus\mathbf{W}_{n-1}^{k}.$ (5)
The operators $H$ and $G$, which transform the basis of the space $V_{j}$ (for $j\in\mathbb{Z}^{+}$) to the bases of the spaces $V_{j-1}$ and $W_{j-1}$ respectively, exhibit the following properties (assuming $H$ and $G$ are real-valued):
$H^{T}H+G^{T}G$ $=I$
$HG^{T}=GH^{T}$ $=0$
$HH^{T}=GG^{T}$ $=I.$
In the two-dimensional discrete wavelet transform, depicted in Fig 4(a), $H$ and $G$ function as low-pass and high-pass filters, respectively, allowing the partitioning of frequency components at each level. Fi 4(b) visually represents this decomposition in a matrix format, enhancing our understanding of how submatrices $B_{n}$, $C_{n}$, and $D_{n}$ are derived from $L_{n}$.
The wavelet transform is defined as $\mathcal{W}:V_{j}\to V_{j-1}\oplus W_{j-1}$ (for $j\in\mathbb{Z}^{+}$), and given by:
$\mathcal{W}=\begin{pmatrix}H\\ G\end{pmatrix}$ (6)
Owing to the orthogonality properties of $H$ and $G$, $\mathcal{W}$ maintains orthogonality. For detailed definitions of $H$, $G$, and $\mathcal{W}$ in 1D and 2D contexts, please refer to Appendix B.
2.3 Similarities between algebraic multigrid and multiresolution
Our objective in comparing these two methodologies is to explore how each enhances the other, as detailed in the following points:
  1. 1.
    The highest resolution space $V_{n}$ in Multiresolution Analysis (MRA) aligns with $\Omega^{h}$, which represents the space of the finest grid vectors in the multigrid method.
  2. 2.
    The decomposition step in MRA, where data is broken down, corresponds to the restriction operation in multigrid methods. Similarly, the forward wavelet transform aligns with the restriction operator $I_{h}^{2h}$.
  3. 3.
    The reconstruction step in MRA, which reassembles the data, is equivalent to the prolongation operation in multigrid methods. Correspondingly, the inverse wavelet transform is equivalent to the prolongation operator $I_{2h}^{h}$.
  4. 4.
    The orthogonality condition $\mathcal{W}\mathcal{W}^{T}=I$ corresponds to the identity relationship for the grid transfer operators, $I_{h}^{2h}I_{2h}^{h}=I$.
  5. 5.
    The multiresolution structure $V_{n}=V_{n-1}\oplus W_{n-1}$ in MRA parallels the multigrid structure $\Omega^{h}=\text{Range}\{I_{2h}^{h}\}\oplus\text{Nullspace}\{I_{h}^{2h}\}$.
These comparisons highlight the synergy between the algebraic multigrid and multiresolution properties of wavelet transforms and the error smoothing and correcting capabilities of intergrid operators in multigrid methods.
Figure 1: Illustration of the multiwavelet-based multigrid method for PDEs. The diagram outlines the method’s workflow, starting with the initial function $f_{n}$ and guess $u_{n}^{0}$. It includes pre-smoothing, multiwavelet-based restriction, solution of correction equations, prolongation back to finer grids, and post-smoothing, illustrating the data flow and transformations within the multigrid process.
3 Multiwavelet-based Algebraic Multigrid Neural Operator
Building on the synergistic qualities of algebraic multigrid and multiresolution wavelet transforms discussed previously, this section elaborates on the Multiwavelet-based Multigrid Neural Operator (M2NO). We begin by defining the problem and its theoretical basis in section 3.1, then detail the construction and practical implementation of multiwavelet-based operators in section 3.2, and finally, illustrate the diverse applications of these methodologies in solving discrete PDEs in section 3.3.
3.1 Problem definition
Formally, let $\mathcal{A}$ and $\mathcal{U}$ denote two Sobolev spaces $\mathcal{H}^{s,p}$ with $s>0$ and $p\geq 1$. In this context, the operator $T$ maps from $\mathcal{A}$ to $\mathcal{U}$, $T:\mathcal{A}\to\mathcal{U}$. We focus on cases where $s>0$ and $p=2$, choosing $p=2$ specifically to enable the definition of projections within a Hilbert space structure, using measures $\mu$.
We define the operator $T$ as an integral operator with a kernel $K$, where $K:D\times D\to L^{2}$. The action of $T$ on an element $a\in\mathcal{A}$ is given by the integral equation:
[…]
This setup allows $T$ to operate within the framework of Hilbert spaces, leveraging the properties of $L^{2}$ spaces to facilitate analysis and computations involving the operator.
An illustrative application of the theory described is the learning of the solution operator for a parametric partial differential equation (PDE). Let $\mathcal{L}_{a}$ denote a differential operator that varies with a parameter $a\in\mathcal{A}$. Consider the general PDE defined as:
$(\mathcal{L}_{a}u)(x)=f(x),\quad x\in D,$ (8)
where $D$ is a bounded, open subset of $\mathbb{R}^{d}$, $f$ is a function residing in an appropriate function space, and the equation is subject to a boundary condition on $\partial D$. Assuming the PDE is well-posed, which implies there exists a unique solution under reasonable assumptions on the data and parameters, we define the operator $F^{\dagger}:\mathcal{A}\to\mathcal{U}$. This operator maps the parameter $a$ to the solution $u$ of (8), effectively encapsulating the dependency of the solution on the parameter within a functional mapping framework.
3.2 Multiwavelet-based Restriction and Prolongation Operators
Given an orthonormal polynomial (OP) basis for $V_{0}^{k}$ as $\{\phi_{0},\phi_{1},\ldots,\phi_{k-1}\}$ with respect to the measure $\mu_{0}$, we can generate the basis for subsequent spaces $V_{n}^{k}$ for $n>1$ through shift and scale (hence the multi-scale aspect) operations on the original basis:
$\phi_{jl}^{n}=2^{\frac{n}{2}}\phi_{j}(2^{n}x-l),\quad j=0,\ldots,k-1,\quad l=0,\ldots,2^{n}-1,\quad\text{w.r.t.}\quad\mu_{n},$ (9)
where $\mu_{n}$ is derived from the shift and scale transformations of $\mu_{0}$. The filter coefficients construct the scaling function $\phi$ and wavelet $\psi$; in our context, functions $\phi$ and $\psi$ are known, and we utilize them to determine the filter coefficients $(H^{(0)},H^{(1)},G^{(0)},G^{(1)})$, as they transform from a fine scale $n=1$ to a coarser scale $n=0$. A uniform measure $\mu_{0}$ version is discussed in [1], and [12] extends it to arbitrary measures by incorporating correction terms $\sum^{(0)}$ and $\sum^{(1)}$.
In one dimension, the operators $H$ and $G$ are constructed as
$H=\begin{pmatrix}H^{(0)}&H^{(1)}\end{pmatrix},\quad G=\begin{pmatrix}G^{(0)}&G^{(1)}\end{pmatrix},$ (10)
to transform the basis of space $V_{j}$ to the bases of spaces $V_{j+1}$ and $W_{j+1}$, respectively. In two dimensions, we utilize the tensor product of one-dimensional multiresolution analyses:
$H=H_{y}\otimes H_{x},\quad G=\begin{pmatrix}G_{y}\otimes H_{x}\\ H_{y}\otimes G_{x}\\ G_{y}\otimes G_{x}\end{pmatrix},$ (11)
where $\otimes$ denotes the Kronecker product. For detailed explanation, see Appendix B.
3.2.1 Low-Pass Filter based Operator
One intuitive way is using the low-pass filter as a restriction operator for the multigrid. The restriction operator and prolongation operator is define as
$I_{h}^{2h}=H,\quad I_{2h}^{h}=(I_{h}^{2h})^{T}=H^{T}.$ (12)
At each decomposition level, it is predominantly the low-pass filters that encapsulate the approximation of the original matrix as illustrated in each one of the levels of scale in Fig 4. In the majority of cases, this methodology proves functional because the residual error associated with the matrices $B$, $C$, and $D$ is almost negligible. However, this property is not maintained by the multigrid method in certain complex scenarios, such as those involving oscillatory coefficients.
Parametrization. With the explicit computation of the low-pass filter $H$, we take the LPF restriction operator as $I_{h}^{2h}=H$ and the corresponding LPF prolongation operator as $I_{2h}^{h}=H^{T}$. It is noteworthy that the prolongation operator is identical for both spaces $V_{0}$ and $V_{1}$, which correspond to the coarsest and the next finer grids, respectively.
3.2.2 Schur complement based Operator
To circumvent these limitations, we introduce a novel set of restriction and prolongation operators, which are founded upon wavelet transformation and the Schur complement. Applying the wavelet transform to both sides of the equation and using the orthogonality of $\mathcal{W}$, we obtain
$(\mathcal{W}\mathcal{L}_{j}\mathcal{W}^{T})\mathcal{W}u=\mathcal{W}f.$ (13)
Denote ~$\tilde{\mathcal{L}}_{j}=\mathcal{W}\mathcal{L}_{j}\mathcal{W}^{T}$, which is partitioned as
~$\tilde{\mathcal{L}}_{j}=\begin{pmatrix}T_{j}&B_{j}\\ C_{j}&D_{j}\end{pmatrix}.$ (14)
Using the block UDL decomposition of ~$\tilde{\mathcal{L}}_{j}$ and the Schur complement [11, 10], the inverse ~$\tilde{\mathcal{L}}_{j}^{-1}$ is expressed as
$u=\begin{pmatrix}H^{T}-G^{T}D_{j}^{-1}C_{j}&G^{T}\end{pmatrix}\begin{pmatrix}(T_{j}-B_{j}D_{j}^{-1}C_{j})^{-1}&0\\ 0&D_{j}^{-1}\end{pmatrix}\begin{pmatrix}H-B_{j}D_{j}^{-1}G\\ G\end{pmatrix}f.$ (15)
The restriction and prolongation operator is defined as
$I_{2h}^{h}=\sqrt{2}(H^{T}-G^{T}D_{j}^{-1}C_{j}),\quad I_{h}^{2h}=\frac{1}{\sqrt{2}}(H-B_{j}D_{j}^{-1}G).$ (16)
For detailed explanations and derivations, please refer to the Appendix B.4.
Parametrization. The multiwavelet-based restriction operator and prolongation operator are approximated using three single-layered CNNs, denoted by $B$ and $C$ and $D^{-1}$, while $H$ and $G$ are precomputed. For a detailed implementation of the model configuration, see Appendix F.2.
3.3 Multiwavelet-based Algebraic Multigrid Method for Discrete PDEs
Employing a linear Finite Element Method (FEM) discretization with mesh size $h=\frac{1}{d}$, the discretized system is described by:
$A_{j}*u_{j}=f_{j},$ (17)
where $u_{j},f_{j}\in\mathbb{R}^{d\times d}$. In this model, * denotes the convolution operation for a single channel, using a $3\times 3$ kernel $A$ determined by the elliptic operator and specific boundary conditions. The inverse of $A*$ aligns with the discrete Green’s function in a linear FEM context. The V-cycle multigrid approach, which can be represented as a conventional convolutional neural network, effectively solves this equation [16, 14].
We provide a concise overview of the essential components of the multigrid structure as shown in Fig 1, framed in the language of convolution as an operator mapping from $f$ to $u$:
  1. 1.
    Initialization: Set the initial function $f_{n}$ as $a_{n}$ and the initial guess $u_{j}^{0}(j=0,\ldots,n)$.
  2. 2.
    Pre-smoothing: Apply $s_{n}$ steps of the smoothing iteration $S_{n}$ to obtain an approximation:
    $u_{n}^{i}=u_{n}^{i-1}+S_{n}^{i}*(f_{n}-A_{n}*u_{n}^{i-1}),\quad i=1,\ldots,s_{n}.$ (18)
  3. 3.
    Multigrid correction: Transition to a coarser grid, where a coarse grid operator $A_{j-1}$ (for $j=1,\ldots,n$) is defined. The residual $r_{j}=f_{j}-A_{j}*u_{j}^{s_{j}}$ is restricted using a multiwavelet-based restriction operator $I_{h}^{2h}$, defined in (16). The correction equation $A_{j-1}e_{j-1}=r_{j-1}$ is then solved over a few relaxation steps $s_{j-1}$. The result, $e_{j-1}$, is prolonged back to the finer grid using the multiwavelet-based prolongation operator $I_{2h}^{h}$, as defined in (16), to update the approximation:
    $f_{j-1}$ $=I_{h}^{2h}(f_{j}-A_{j}*u_{j}^{s_{j}}),$ (19)
    $u_{j-1}^{i}$ $=u_{j-1}^{i-1}+S_{j-1}^{i}*f_{j-1},\quad i=1,\ldots,s_{j-1},$ (20)
    $\tilde{u}_{j}$ $=u_{j}^{s_{j}}+I_{2h}^{h}u_{j-1}^{s_{j-1}}.$ (21)
  4. 4.
    Post-processing: At this stage, one may choose to further prolongate to a finer level in a process known as super-resolution or continue with post-smoothing steps.
The following theorem establishes the convergence properties of the V-cycle multigrid operator. The detailed proof of this theorem is provided in the Appendix C.
Theorem 1 (Operator Convergence [34]).
Let us denote the V-cycle multigrid operator as $\mathcal{L}_{mg}$. The convergence can be succinctly expressed by the inequality
$\lVert u-\mathcal{L}_{mg}(A*u)\rVert_{A}\leq\left(1-\frac{1}{c}\right)\lVert u\rVert_{A},$ (22)
where $c$ represents a constant that remains unaffected by the discretization parameter $h$.
The above theorem affirms the uniform approximation efficacy of $\mathcal{L}_{mg}$ in relation to the pseudo-inverse of $A*$, associated with the Green’s function for the elliptic operator $L_{n}$. The notation $\lVert u\rVert_{A}^{2}=(u,A*u)_{\mathbb{L}^{2}(\Omega)}$ defines the energy norm within the domain $\Omega$.
Table 1: Performance comparison with baselines on benchmarks. $L_{2}$ loss is recorded. For clarity, the best result is in bold and the second best is underlined. The results presented in this table were obtained at the highest resolution settings.
Model 1D 2D
Advection Burgers Darcy Navier-Stoke Diffusion-Reaction
$\nu=1e-3$ $\beta=4.0$ $\nu=1e-3$ $\nu=1e-5$
UNet 3.44E-05 3.83E-05 1.54E-01 3.86E-03 2.15E-02 1.46E-01
FNO-2D 1.68E-05 1.81E-05 1.95E-02 4.95E-03 1.72E-02 1.64E-01
MGNO / / 5.72E-02 3.39E-03 1.27E-02 1.12E-01
[…]
LSM / / 6.69E-02 6.00E-03 1.71E-02 1.58E-01
M2NO 2.14E-06 7.16E-06 2.73E-02 9.17E-04 4.07E-03 1.01E-01
[…]
Promotion 78.41% 60.49% 49.56% 49.75% 65.77% 10.15%
4 Experiments
In this section, we evaluate the performance of the M2NO model on a variety of PDE benchmarks, incorporating different boundary conditions, and conduct Multiresolution Analysis. Specifically, we focus on two variants: M2NO with a Schur complement based Operator and M2NO-LPF featuring a Low-Pass Filter based Operator. These variants are assessed to understand their respective impacts on solving complex PDEs under varied computational scenarios.
General Setting. We consider a set of function pairs $(a_{i},u_{i})_{i=1}^{N}$, where $a_{i}=a(x_{i})$ and $u_{i}=u(x_{i})$ at discretized points $x_{i}$ in the domain $D$. We express the functions $f_{a}$ and $f_{u}$ as linear combinations of basis functions $\phi_{ji}^{n}$, formulated as $f_{a}=\sum_{i=1}^{N}a_{i}\phi_{ji}^{n}$ and $f_{u}=\sum_{i=1}^{N}u_{i}\phi_{ji}^{n}$, both of which belong to the space $V_{n}^{k}$ where $n=\log_{2}N$. These formulations serve as notational conventions rather than explicit constructs. Given the dataset $(a_{i},u_{i})_{i=1}^{N}$, our goal is to approximate the operator $\mathcal{G}$ by optimizing network parameters $\theta$. This involves solving the following optimization problem:
$\min_{\theta\in\Theta}\mathcal{L}(\theta):=\min_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\left[\lVert\tilde{\mathcal{G}}_{\theta}(a_{i})-u_{i}\rVert^{2}\right].$ (23)
Benchmarks. To substantiate the superiority of our method, we conducted comprehensive method comparisons across multiple benchmark scenarios [22, 30]. For clearness, we summarize the benchmarks in Table 4 and show the examples of PDEs in Fig 5. More details about PDE benchmarks can be found in Appendix D.
Baselines. We conducted a comprehensive comparison against established methods including UNet [26], Fourier Neural Operator (FNO) [22], Multiwavelet Transform (MWT) [12], Multigrid Neural Operator (MgNO) [15], and Latent Spectral Model (LSM) [33]. It is important to note that MgNO and LSM are specifically tailored for 2D problems.
Implementation. For fairness, all methods were trained over 500 epochs using the $L_{2}$ loss function [22], with the ADAM optimizer [18] starting at a learning rate of $10^{-3}$ and halving every 100 epochs ($\gamma=0.5$). Batch sizes were set at 512 for 1D PDEs and 16 for 2D PDEs. Unless specified, baselines used default hyperparameters. Experiments were performed on a single Nvidia A800 80GB GPU.
For further details on the baseline models and experimental setup, refer to Appendix F.1.
4.1 Main Results
Table 1 presents a performance comparison across five PDE benchmarks, highlighting M2NO’s superior accuracy. For the 1D benchmarks, M2NO achieves $L_{2}$ losses of 2.14E-06 for the Advection equation and 7.16E-06 for the Burgers equation, demonstrating improvements of 78.41% and 60.49% respectively over the next best model. In the 2D benchmarks, M2NO consistently outperforms other models, achieving the lowest $L_{2}$ losses of 9.17E-04, 4.07E-03 and 1.01E-01 for the Navier-Stokes with $\nu=1e-3$, $\nu=1e-5$ and Diffusion-Reaction equations respectively, marking significant enhancements of 49.75%, 65.77% and 10.15% respectively compared to the next best results. These findings not only underscore the exceptional accuracy of M2NO but also its robust performance across various challenging computational settings.
Figure 2: Performance showcases on the Navier-Stokes benchmark, including plots of prediction errors for enhanced clarity.
M2NO’s Superiority in Handling High-Frequency Dynamics. As demonstrated in Fig 2, M2NO excels at mitigating high-frequency errors across various models, particularly within dynamic environments such as the Navier-Stokes benchmark. While the M2NO-LPF variant, equipped with a low-pass filter, outperforms M2NO in time-independent tasks like the Darcy dataset by better maintaining global structures, M2NO consistently surpasses other models in scenarios with rapid, localized changes. This ability to effectively handle high-frequency information underscores M2NO’s advantage in time-dependent datasets, highlighting its suitability for complex, dynamic simulations.
Benchmarking M2NO on Complex PDEs. The 2D diffusion-reaction equation is chosen as a benchmark due to its complexity, involving two key variables—the activator and the inhibitor—which are non-linearly coupled. This non-linear coupling presents a significant challenge in accurately modeling the dynamic interactions within the system. Demonstrating its robust modeling capabilities, M2NO secured an improvement exceeding 10% on this challenging benchmark, underscoring its proficiency in handling intricate partial differential equations.
For further experimental results, including outcomes for each PDE under various hyperparameters, please refer to Appendix D. For ablation and hyperparameter study, refer to Appendix E.
4.2 Multiresolution Analysis
[…]
Figure 3: Model performance on the Burgers benchmark at resolutions ranging from 128 to 8192.
Performance under Various Resolutions. As illustrated in Fig. 3, the M2NO model consistently outperforms baseline methods across all tested resolutions in the high-resolution Burgers dataset [22]. This demonstrates its robust capacity to handle the complexities of high-dimensional PDEs without the need for parameter readjustment, even as resolution increases. The M2NO model’s superior performance is further highlighted by its ability to efficiently leverage its multiresolution framework to extract and process information effectively across a wide range of inputs. Detailed results across various datasets and resolutions can be found in Appendix D, offering a comprehensive view of M2NO’s capabilities in diverse computational scenarios.
Table 2: Performance of MWT, M2NO and M2NO-LPF models on the super-resolution task, showing $L_{2}$ losses at various resolutions following training at $s=128$.
Model $s=256$ $s=512$ $s=1024$ $s=2048$ $s=4096$ $s=8192$
MWT 3.17E-02 3.33E-02 3.38E-02 3.68E-02 3.89E-02 4.56E-02
[…]
M2NO-LPF 3.25E-03 3.31E-03 2.87E-03 2.76E-03 2.63E-03 2.56E-03
Super-Resolution Performance. Table 2 illustrates the super-resolution capabilities of the MWT, M2NO, and M2NO-LPF models, trained at a base resolution of $s=128$. This evaluation assesses their ability to upscale to higher resolutions, a critical feature of the M2NO framework that utilizes multiwavelet-based multigrid methods. M2NO-LPF consistently records the lowest $L_{2}$ losses across all tested resolutions, demonstrating its superior ability to capture fine details essential for high-resolution tasks. These results highlight the effectiveness of integrating multiwavelet transformations with a multigrid approach in enhancing super-resolution prediction accuracy.
5 Conclusion
We introduced the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), a novel framework for solving high-dimensional (referring to the increase in the number of grid points or resolution) partial differential equations using a synergistic blend of multiwavelet transformations and algebraic multigrid techniques. M2NO excels in capturing both global trends and detailed features within PDE solutions, demonstrating superior accuracy and scalability across diverse datasets.
While the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO) is particularly tailored for complex, time-dependent PDEs with high-frequency and localized changes, this specialization represents a significant advantage, as most practical PDEs exhibit these characteristics. Its current limitation in handling non-uniform grids, which can be adapted through transformations, presents an opportunity for direct integration to further enhance efficiency. Moving forward, efforts will focus on broadening M2NO’s utility across a more diverse range of PDE challenges and simplifying its deployment to maximize its impact in scientific computing.
Appendix A Algebraic Multigrid Methods
The key ingredients are the operators $I_{h}^{2h}$ and $I_{2h}^{h}$ that change grids:
  • A restriction operator $I_{h}^{2h}$ transfers vectors from the fine grid to the coarse grid
  • An interpolation operator $I_{2h}^{h}$ returns to the fine grid
  • The original matrix $A_{h}$ on the fine grid is approximated by $A_{2h}=I_{h}^{2h}A_{h}I$ on the coarse grid.
The spectral radius $\rho(I_{h}^{2h})$ is called convergence factor. It is roughly the worst factor by which the error is reduced with each relaxation sweep.
A.1 Some Theory
Recall the variational properties:
[…]
$I^{2h}_{h}$ $=c(I^{h}_{2h})^{T},$ $c\in I_{h}^{2h}.$
From the orthogonality relationships between the subspaces of a linear operator (Fundamental Theorem of Linear Algebra), we know that
$\text{Nullspace}(I^{2h}_{h})\;\bot\;\text{Range}[(I^{2h}_{h})^{T}].$ (25)
By the second variational property, it then follows that:
$\text{Nullspace}(I^{2h}_{h})\;\bot\;\text{Range}(I^{h}_{2h}),$ (26)
which means that $(\mathbf{q}^{h},\mathbf{u}^{h})=0$ whenever $\mathbf{q}^{h}\in\text{Range}(I_{2h}^{h})$ and $I_{h}^{2h}\mathbf{u}^{h}=0$. This is equivalent to the condition that $(\mathbf{q}^{h},A^{h}\mathbf{u}^{h})=0$ whenever $\mathbf{q}^{h}\in\text{Range}(I_{2h}^{h})$ and $I_{h}^{2h}A^{h}\mathbf{u}^{h}=0$.
This orthogonality property allows the space $\Omega^{h}$ to be decomposed in the form:
$\Omega^{h}=\text{Range}(I_{2h}^{h})\oplus\text{Nullspace}(I_{h}^{2h}A^{h}),$ (27)
which means that if $\mathbf{e}^{h}$ is a vector in $\Omega^{h}$, then it may always be expressed as
$\mathbf{e}^{h}=\mathbf{s}^{h}+\mathbf{t}^{h},$ (28)
where $\mathbf{s}^{h}\in\text{Range}(I_{2h}^{h})$ and $\mathbf{t}^{h}\in\text{Nullspace}(I_{h}^{2h}A^{h})$. We associate $\mathbf{s}^{h}$ with the smooth components of the error and we associate $\mathbf{t}^{h}$ with the oscillatory components of the error.
Algorithm 1 V-cycle
1:procedure Vcycle($v^{h}$, $f^{h}$)
2:     Relax $\alpha_{1}$ times on $A^{h}u^{h}=f^{h}$ with initial guess $v^{h}$
3:     if $\Omega^{h}$ is the coarsest grid then
4:         Relax $\alpha_{2}$ times on $A^{h}u^{h}=f^{h}$ with initial guess $v^{h}$
5:     else
6:         $r^{h}\leftarrow f^{h}-A^{h}v^{h}$triangleright Compute the residual
7:         $f^{2h}\leftarrow I^{2h}_{h}r^{h}$triangleright Restrict the residual to the coarser grid
8:         $v^{2h}\leftarrow 0$triangleright Initial guess for the coarser grid
9:         $v^{2h}\leftarrow$ Vcycle($v^{2h},f^{2h}$) $\triangleright$ Recursive call
10:         $v^{h}\leftarrow v^{h}+I^{h}_{2h}v^{2h}$triangleright Prolongate the correction and update
11:         Relax $\alpha_{2}$ times on $A^{h}u^{h}=f^{h}$ with initial guess $v^{h}$
12:     end if
13:end procedure
Appendix B Multiwavelets
B.1 Multiresolution analysis
For $k=1,2,\ldots$, and $n=0,1,2,\ldots$, we define $\mathbf{V}_{n}^{k}$ as a space of piecewise polynomial functions,
[…]
$\text{of degree less than }k\text{, for }l=0,\ldots,2^{n}-1\text{, and }f\text{ vanishes elsewhere}\}.$
The space $\mathbf{V}_{n}^{k}$ has dimension $2^{n}k$ and
$\mathbf{V}_{0}^{k}\subset\mathbf{V}_{1}^{k}\subset\ldots\subset\mathbf{V}_{n}^{k}\subset\ldots.$ (30)
We define the multiwavelet subspace $\mathbf{W}_{n}^{k}$, $n=0,1,2,\ldots$, as the orthogonal complement of $\mathbf{V}_{n}^{k}$ in $\mathbf{V}_{n+1}^{k}$,
$\mathbf{V}_{n}^{k}\oplus\mathbf{W}_{n}^{k}=\mathbf{V}_{n+1}^{k},\quad\mathbf{W}_{n}^{k}\bot\mathbf{V}_{n}^{k},$ (31)
and note that $\mathbf{W}_{n}^{k}$ is of dimension $2^{n}k$. Therefore, we have
$\mathbf{V}_{n}^{k}=\mathbf{V}_{0}^{k}\oplus\mathbf{W}_{0}^{k}\oplus\mathbf{W}_{1}^{k}\oplus\ldots\oplus\mathbf{W}_{n-1}^{k}.$ (32)
We define $\mathbf{V}^{k}=\bigcup\limits_{n=0}^{\infty}{V_{n}^{k}}$ and observe that $\mathbf{V}^{k}$ is dense in $L^{2}([0,1])$ with respect to norm $\|{f}\|=\langle{f,f}\rangle^{1/2}$, where
[…]
Given a basis $\phi_{0},\ldots,\phi_{k-1}$ of $\mathbf{V}_{0}^{k}$, the space $\mathbf{V}_{n}^{k}$ is spanned by $2^{n}k$ functions which are obtained from $\phi_{0},\ldots,\phi_{k-1}$ by dilation and translation,
$\phi_{jl}^{n}=2^{n/2}\phi_{j}(2^{n}x-l),\quad j=0,\ldots,k-1,\quad l=0,\ldots,2^{n}-1.$ (34)
B.2 Multiwavelets
We introduce piecewise polynomial functions $\psi_{0}\ldots,\psi_{k-1}$ to be an orthonormal basis for $\mathbf{W}_{0}^{k}$,
[…]
Since $\mathbf{W}_{0}^{k}\bot\mathbf{V}_{0}^{k}$, the first $k$ moments of $\psi_{0},\ldots,\psi_{k-1}$ vanish:
[…]
The space $\mathbf{W}_{n}^{k}$ is spanned by $2^{n}k$ functions obtained from $\psi_{0},\ldots,\psi_{k-1}$ by dilation and translation,
$\psi_{jl}^{n}(x)=2^{n/2}\psi_{j}(2^{n}x-l),\quad j=0,\ldots,k-1,\quad l=0,\ldots,2^{n}-1,$ (37)
and supp $(\psi_{jl}^{n})=I_{nl}$ denotes the interval $[2^{-n}l,2^{-n}(l+1)]$. The condition of orthonormality of $\psi_{0},\ldots,\psi_{k-1}$ yields
[…]
The set $\{\phi_{0},\ldots,\phi_{k-1}\}\cup\{\psi_{jl}^{n}:j=0,\ldots,k-1,l=0,\ldots,2^{n}-1,n=0,1,\ldots\}$ therefore forms a complete orthonormal basis for $\mathbb{L}^{2}([0,1])$.
B.3 Multiwavelet Transformation
In order to compute projections of functions on subspaces of multiresolution analysis, we consider the two-scale difference equations. In our case, functions $\phi$ and $\psi$ are known, and we use them to construct the filter coefficients. The relations (30) and (31) between the subspaces may be expressed by the two-scale difference equations,
$\phi_{i}(x)=\sqrt{2}\sum_{j=0}^{k-1}{(h_{ij}^{(0)}\phi_{j}(2x)-h_{ij}^{(1)}\phi_{j}(2x-1))},\quad i=0,\ldots,k-1,$ (39)
$\psi_{i}(x)=\sqrt{2}\sum_{j=0}^{k-1}{(g_{ij}^{(0)}\phi_{j}(2x)+g_{ij}^{(1)}\phi_{j}(2x-1))},\quad i=0,\ldots,k-1,$ (40)
where the coefficients $g_{ij}^{(0)}$, $g_{ij}^{(1)}$ depend on the choice of the order $k$. The function $\sqrt{2}\phi_{0}(2x)\ldots,\sqrt{2}\phi_{k-1}(2x)$ in (39) are orthonormal on the interval $[0,\frac{1}{2}]$ whereas $\sqrt{2}\phi_{0}(2x-1)\ldots,\sqrt{2}\phi_{k-1}(2x-1)$ are orthonormal on the interval $[\frac{1}{2},1]$. The matrices of coefficients
$H^{(0)}=\{h_{ij}^{(0)}\},\quad H^{(1)}=\{h_{ij}^{(1)}\},\quad G^{(0)}=\{g_{ij}^{(0)}\},\quad G^{(1)}=\{g_{ij}^{(1)}\}$ (41)
are analogs of the quadrature mirror filters. The two-scale equations (39),(40) lead us to a multiresolution decomposition.
By construction, we have $\langle{\phi_{i},\phi_{j}}\rangle=\delta_{ij}$, $\langle{\psi_{i},\psi_{j}}\rangle=\delta_{ij}$, and $\langle{\phi_{i},\phi_{j}}\rangle=0$ for $i,j=0,\ldots,k-1$, where $\langle$,$\rangle$ , ⟩ is the inner product defined in (33). Using this orthogonality conditions and (39-40), we obtain
$H^{(0)}H^{(0)T}+H^{(1)}H^{(1)T}$ $=I,$ (42)
$G^{(0)}G^{(0)T}+G^{(1)}G^{(1)T}$ $=I,$ (43)
$H^{(0)}G^{(0)T}+H^{(1)}G^{(1)T}$ $=0.$ (44)
The matrices of coefficients $H^{(0)},H^{(1)},G^{(0)},G^{(1)}$ are analogs of the quadrature mirror filter. We now derive the necessary relations for multiresolution reconstruction. In one dimension, we construct
$H$ $=\begin{pmatrix}H^{(0)}&H^{(1)}\end{pmatrix}$ (45)
$G$ $=\begin{pmatrix}G^{(0)}&G^{(1)}\end{pmatrix}$
to be the operators that transform the basis of the space $V_{j}$ to the bases of the spaces $V_{j+1}$ and $W_{j+1}$, respectively. And $H$ and $G$ play the role of separating the low and high frequency components. With the property of $H^{(0)},H^{(1)},G^{(0)},G^{(1)}$, we have
$H^{T}H+G^{T}G$ $=I$ (46)
$HG^{T}=GH^{T}$ $=0$
$HH^{T}=GG^{T}$ $=I$
In two dimensions, the tensor product of one-dimensional multiresolution analyses is used. Analogous to the one-dimensional case, define the operators $H$ and $G$ so that
$H$ $=H_{y}\otimes H_{x}$ (47)
$G$ $=\begin{pmatrix}G_{y}\otimes H_{x}\\ H_{y}\otimes G_{x}\\ G_{y}\otimes G_{x}\end{pmatrix},$
where $\otimes$ is the Kronecker product.
We denote the multiwavelet transform as
$\mathcal{W}=\begin{pmatrix}H\\ G\end{pmatrix},$ (48)
Due to the properties of $H$ and $G$ , $\mathcal{W}$ is orthogonal.
B.4 Multiwavelet-based Multigrid method
The following is a brief description of the two-dimensional application of the multiwavelet multigrid method. Given the problem
$\mathcal{L}_{j}u=f,$ (49)
where $\mathcal{L}_{j}$ represents the operator obtained by discretizing a two-dimensional partial differential equation on the fine grid, apply the wavelet transform to both sides of the equation and use the orthogonality of $\mathcal{W}_{j}$ to obtain
$(\mathcal{W}_{j}\mathcal{L}_{j}\mathcal{W}_{j}^{T})\mathcal{W}_{j}u=\mathcal{W}_{j}f.$ (50)
Denote ~$\tilde{\mathcal{L}}_{j}=\mathcal{W}_{j}\mathcal{L}_{j}\mathcal{W}_{j}^{T}$, is partitioned to obtain
~$\tilde{\mathcal{L}}_{j}=\mathcal{W}_{j}\mathcal{L}_{j}\mathcal{W}_{j}^{T}=\begin{pmatrix}T_{j}&B_{j}\\ C_{j}&D_{j}\end{pmatrix}.$ (51)
Using the block UDL decomposition of ~$\tilde{\mathcal{L}}_{j}$ and the Schur complement of inverse ~$\tilde{\mathcal{L}}_{j}^{-1}$, we can get
$u=\begin{pmatrix}H^{T}-G^{T}D_{j}^{-1}C_{j}&G^{T}\end{pmatrix}\begin{pmatrix}(T_{j}-B_{j}D_{j}^{-1}C_{j})^{-1}&0\\ 0&D_{j}^{-1}\end{pmatrix}\begin{pmatrix}H-B_{j}D_{j}^{-1}G\\ G\end{pmatrix}f.$ (52)
Denote the prolongation operator
$I_{2h}^{h}=\sqrt{2}(H^{T}-G^{T}D_{j}^{-1}C_{j}),$ (53)
as well as the restriction operator
$I_{h}^{2h}=\frac{1}{\sqrt{2}}(H-B_{j}D_{j}^{-1}G).$ (54)
Plugging the prolongation and restriction operators into (15) gives
$e=I_{2h}^{h}(T_{j}-B_{j}D_{j}^{-1}C_{j})^{-1}I_{h}^{2h}r+G_{j}^{T}D_{j}^{-1}G_{j}r.$ (55)
Assuming that $G_{j}^{T}D_{j}^{-1}G_{j}r$ is small, i.e., $r$ is almost in $Range(H_{j}^{T})$, the error can be approximated by
$e=I_{2h}^{h}(T_{j}-B_{j}D_{j}^{-1}C_{j})^{-1}I_{h}^{2h}r,$ (56)
resulting in
$(T_{j}-B_{j}D_{j}^{-1}C_{j})e_{2h}=I_{h}^{2h}r_{h}.$ (57)
By substituting the prolongation operator and restriction operator defined by multiwavelets into the V-cycle Algorithm 1, the Multiwavelet Multigrid method can be obtained.
(a) The two-dimensional discrete wavelet transform.
(b) Matrix representation of decomposition.
Figure 4: (a) Two-dimensional discrete wavelet transform with $G$ (high-pass) and $H$ (low-pass) filters. Horizontal lines represent frequency bands; vertical lines show decomposition levels. (b) Matrix representation of decomposition, dividing $L_{n}$ into submatrices $B_{n}$, $C_{n}$, and $D_{n}$.
Appendix C Proof of Theorem 1
The proof relies on several key assumptions and properties of the operators involved:
  1. 1.
    Assumptions on Subspace Solvers: We assume that the subspace solvers $T_{i}$ satisfy the following conditions:
    • $R(T_{i})=V_{i}$ and $T_{i}:V_{i}\to V_{i}$ is isomorphic for each $i$.
    • Each factor $I-T_{i}$ is nonexpansive, i.e., $\|I-T_{i}\|\leq 1$, which is equivalent to $(T_{i}v,T_{i}v)\leq 2(T_{i}v,v)$ for all $v\in V$.
  2. 2.
    Symmetrization of Operators: The symmetrized operator is given by $\bar{T}_{i}=T_{i}+T_{i}^{*}-T_{i}^{*}T_{i}$. This operator is nonnegative on $V$ and symmetric positive definite on $V_{i}$.
  3. 3.
    Main Result: Under the above assumptions, we establish that:
    $\|E\|_{L(V,V)}\leq\left(1+\sup_{\|v\|=1}\inf_{\sum v_{i}=v}\sum_{i=1}^{J}(\bar{T}_{i}^{-1}T_{i}^{*}w_{i},T_{i}^{*}w_{i})\right)^{-1}.$ (58)
Table 3: Definitions of Symbols.
Symbol Description
$\mathcal{L}_{mg}$ V-cycle multigrid operator
$A$ Elliptic operator
$\ast$ Convolution operator
$\|\cdot\|_{A}$ Energy norm defined as $\|u\|_{A}^{2}=(u,A\ast u)_{L^{2}(\Omega)}$
$c$ A constant unaffected by the discretization parameter $h$
$\Omega$ Domain of the problem
$T_{i}$ Subspace solvers
$V_{i}$ Subspaces
$I$ Identity operator
$\bar{T}_{i}$ Symmetrized operator $\bar{T}_{i}=T_{i}+T_{i}^{*}-T_{i}^{*}T_{i}$
$L_{n}$ Elliptic operator associated with the Green’s function
$w_{i}$ Intermediate vectors defined by $w_{i}=\sum_{j=i}^{J}v_{j}-T_{i}^{-1}v_{i}$
$\lambda_{\min}(\bar{T}_{i})$ Smallest eigenvalue of $\bar{T}_{i}$
To complete the proof, we need to show that the constant $c_{0}$ in the main result is finite. This involves demonstrating that the series involving $T_{i}$ and $\bar{T}_{i}$ is bounded. The definitions of symbols are summarized in Table. 3.
First, we need to show that the constant $c_{0}$ is finite. We define $c_{0}$ as follows:
$c_{0}=\sup_{\|v\|=1}\inf_{\sum v_{i}=v}\sum_{i=1}^{J}(\bar{T}_{i}^{-1}T_{i}^{*}w_{i},T_{i}^{*}w_{i}),$ (59)
where $w_{i}=\sum_{j=i}^{J}v_{j}-T_{i}^{-1}v_{i}$.
Using the properties of the symmetrized operator $\bar{T}_{i}$ and the nonexpansive property of $T_{i}$, we have:
$(T_{i}v,T_{i}v)\leq 2(T_{i}v,v)\leq 2\|T_{i}v\|\|v\|\leq 2\|T_{i}\|\|v\|^{2}.$ (60)
Since $\bar{T}_{i}$ is symmetric positive definite on $V_{i}$, we have:
$(\bar{T}_{i}v,v)\geq\lambda_{\min}(\bar{T}_{i})\|v\|^{2},$ (61)
where $\lambda_{\min}(\bar{T}_{i})$ is the smallest eigenvalue of $\bar{T}_{i}$.
We now show that the series involving $T_{i}$ and $\bar{T}_{i}$ is bounded:
$\sum_{i=1}^{J}(\bar{T}_{i}^{-1}T_{i}^{*}w_{i},T_{i}^{*}w_{i})\leq\sum_{i=1}^{J}\frac{1}{\lambda_{\min}(\bar{T}_{i})}(T_{i}^{*}w_{i},T_{i}^{*}w_{i}).$ (62)
Since $\|T_{i}\|\leq 1$ and $\|T_{i}^{*}\|\leq 1$, we have:
$(T_{i}^{*}w_{i},T_{i}^{*}w_{i})\leq\|T_{i}^{*}\|^{2}\|w_{i}\|^{2}\leq\|w_{i}\|^{2}.$ (63)
Therefore,
$\sum_{i=1}^{J}(\bar{T}_{i}^{-1}T_{i}^{*}w_{i},T_{i}^{*}w_{i})\leq\sum_{i=1}^{J}\frac{1}{\lambda_{\min}(\bar{T}_{i})}\|w_{i}\|^{2}.$ (64)
We need to show that $\|w_{i}\|$ is bounded. By the definition of $w_{i}$, we have:
$w_{i}=\sum_{j=i}^{J}v_{j}-T_{i}^{-1}v_{i}.$ (65)
Using the fact that $T_{i}^{-1}$ is bounded, we get:
$\|w_{i}\|\leq\left\|\sum_{j=i}^{J}v_{j}\right\|+\|T_{i}^{-1}\|\|v_{i}\|\leq\sum_{j=i}^{J}\|v_{j}\|+\|T_{i}^{-1}\|\|v_{i}\|.$ (66)
Since $\|v\|=1$ and $v=\sum_{i=1}^{J}v_{i}$, we have:
$\|v\|^{2}=\left\|\sum_{i=1}^{J}v_{i}\right\|^{2}\leq J\sum_{i=1}^{J}\|v_{i}\|^{2}\leq J.$ (67)
Thus,
$\|w_{i}\|\leq\sum_{j=i}^{J}\|v_{j}\|+\|T_{i}^{-1}\|\|v_{i}\|\leq J+\|T_{i}^{-1}\|.$ (68)
Combining the above results, we get:
$\sum_{i=1}^{J}(\bar{T}_{i}^{-1}T_{i}^{*}w_{i},T_{i}^{*}w_{i})\leq\sum_{i=1}^{J}\frac{1}{\lambda_{\min}(\bar{T}_{i})}(J+\|T_{i}^{-1}\|)^{2}.$ (69)
Since $\lambda_{\min}(\bar{T}_{i})$ and $\|T_{i}^{-1}\|$ are finite, the constant $c_{0}$ is finite.
Finally, combining these results, we have:
$\|u-\mathcal{L}_{mg}(A\ast u)\|_{A}\leq\left(1-\frac{1}{c}\right)\|u\|_{A},$ (70)
where $c$ is the aforementioned constant.
Thus, the operator $\mathcal{L}_{mg}$ demonstrates uniform convergence with the desired properties.
Appendix D Details for Benchmarks and Additional Experiments
Figure 5: Examples of PDE benchmarks across different scenarios. Left side (1D): (a) Advection, where input and output represent the fluid quantity being advected at different times; (b) Burgers, showcasing the evolution of fluid velocity. Right side (2D): (c) Darcy Flow, depicting changes in fluid velocity influenced by diffusion coefficients; (d) Navier-Stokes, illustrating fluid velocity variations over time; (e) Diffusion-Reaction, visualizing the interaction dynamics of activators and inhibitors.
We detail the benchmark configurations in Table 4 and illustrate PDE examples in Figure 5. Below, we outline the generation process for each category of governing PDEs and present comprehensive experimental results for each benchmark, exploring variations across different coefficients and resolutions.
For the Navier-Stokes benchmark, we utilize the dataset released by [22], available for download at this URL with MIT License. For other benchmarks, we employ widely recognized PDEs from the PDE benchmark suite, a popular resource frequently used in contemporary methods. These benchmarks can be accessed at this URL under the CC BY 4.0 License.
Table 4: Summary of benchmarks. Each benchmark varies by dimension, time dependency, boundary conditions, and resolution. Specific parameters include the viscosity ($\nu$) values for the Burgers benchmarks, set at {0.001, 0.01, 0.1, 1.0}, and diffusion coefficients ($\beta$) for the Advection benchmark, set at {0.1, 1.0, 4.0}.
Benchmark # Dimension Time dependent Boundary condition Spatial resolution Temporal resolution
Burgers 1D Yes Periodic 1024 200
Burgers-HR 1D Yes Periodic 8192 -
Advection 1D Yes Periodic 1024 200
Darcy Flow 2D No Dirichlet 512*512 -
Navier Stokes 2D Yes Periodic 64*64 20
Diffusion-Reactiong 2D Yes Periodic 128*128 100
D.1 1D Advection Equation
The advection equation [30] models pure advection behavior without non-linearity. Its expression is given as:
$\frac{\partial u(t,x)}{\partial t}+\beta\frac{\partial u(t,x)}{\partial x}$ $=0,\quad x\in(0,1),\ t\in(0,2],$ (71)
$u(0,x)$ $=u_{0}(x),\quad x\in(0,1),$
where $\beta$ is a constant advection speed.
As the initial condition, we use a superposition of sinusoidal waves:
$u_{0}(x)=\sum_{k_{i}=k_{1}}^{k_{N}}A_{i}\sin(k_{i}x+\phi_{i}),$ (72)
where $k_{i}=\frac{2\pi n_{i}}{L_{x}}$ are wave numbers with $n_{i}$ being integers randomly selected in $[1,n_{\text{max}}]$, $A_{i}$ are random float numbers uniformly chosen in $[0,1]$, and $\phi_{i}$ are phases randomly chosen in $(0,2\pi)$. We set $k_{\text{max}}=8$ and $N=2$.
The exact solution of the system is given by:
$u(t,x)=u_{0}(x-\beta t).$ (73)
In our dataset, we consider only periodic boundary conditions.
Table 5: Benchmarks on 1D Advection equation at various $\beta$ and input resolution s.
Model beta=0.1 beta=1.0 beta=4.0
s=128 s=256 s=512 s=1024 s=128 s=256 s=512 s=1024 s=128 s=256 s=512 s=1024
Unet 3.32E-05 2.42E-05 3.21E-05 2.94E-05 3.05E-05 1.54E-05 3.52E-05 3.52E-05 3.79E-05 2.59E-05 4.32E-05 3.44E-05
FNO 1.11E-05 5.79E-06 4.58E-06 6.27E-06 2.19E-05 2.16E-05 2.74E-05 2.17E-05 2.09E-05 2.10E-05 1.70E-05 1.68E-05
MGNO / / / / / / / / / / / /
MWT Leg 8.74E-06 3.44E-06 8.55E-06 8.14E-06 3.12E-06 4.06E-06 4.30E-06 4.41E-06 3.50E-06 6.55E-06 9.89E-06 9.91E-06
LSM / / / / / / / / / / / /
[…]
Promotion 89.24% 76.16% 87.12% 91.23% 55.13% 75.37% 76.51% 75.96% 59.71% 66.56% 76.34% 78.41%
Results. Table 5 highlights the M2NO model’s superior performance in solving the 1D Advection equation across various diffusion coefficients ($\beta$) and resolutions ($s$). M2NO consistently outperforms traditional models like U-net and FNO, particularly at higher $\beta$ values, which suggest increased diffusivity. Notably, M2NO-LPF performs well at lower resolutions, indicating its effectiveness in capturing essential features with limited data. These results demonstrate M2NO’s robust adaptability to different physical and computational conditions, reinforcing its potential for complex PDE challenges.
D.2 1D Burgers’ Equation
The Burgers’ equation [30, 22, 12] is a PDE that models non-linear behavior and diffusion processes in fluid dynamics. It is expressed as:
$\frac{\partial u}{\partial t}$ $=-u\frac{\partial u}{\partial x}+\nu\frac{\partial^{2}u}{\partial x^{2}},\quad x\in(0,1),\ t\in(0,2],$ (74)
$u(0,x)$ $=u_{0}(x),\quad x\in(0,1),$
where $\nu$ is the diffusion coefficient, which is assumed to be constant in our dataset.
Table 6: Benchmarks on 1D Burgers equation at various $\nu$ and input resolution s.
Model nu=1.0 nu=0.1 nu=0.01 nu=0.001
s=128 s=256 s=512 s=1024 s=128 s=256 s=512 s=1024 s=128 s=256 s=512 s=1024 s=128 s=256 s=512 s=1024
Unet 8.27E-05 4.78E-05 3.16E-05 6.10E-05 4.33E-05 4.26E-05 4.56E-05 3.74E-05 5.30E-05 4.98E-05 2.59E-05 4.99E-05 8.29E-05 6.66E-05 6.04E-05 3.83E-05
FNO 9.18E-06 1.15E-05 2.71E-05 3.14E-05 8.00E-06 8.40E-06 8.70E-06 8.06E-06 1.70E-05 1.21E-05 1.10E-05 1.10E-05 4.65E-05 3.48E-05 2.28E-05 1.81E-05
MGNO / / / / / / / / / / / / / / / /
MWT Leg 3.46E-06 4.66E-06 5.27E-06 6.49E-06 3.84E-06 3.48E-06 3.48E-06 3.63E-06 1.81E-05 1.17E-05 1.15E-05 1.18E-05 5.24E-05 3.69E-05 2.50E-05 2.07E-05
LSM / / / / / / / / / / / / / / / /
[…]
Promotion 54.91% 67.60% 55.22% 59.01% 78.39% 64.94% 54.60% 49.59% 81.24% 71.95% 74.77% 67.40% 51.96% 73.07% 58.26% 60.49%
Results. Table 6 showcases the performance of models on the 1D Burgers equation across different viscosities ($\nu$) and resolutions ($s$). The M2NO model consistently demonstrates low $L_{2}$ losses, particularly excelling at higher viscosities where the equation dynamics are less complex. Conversely, at lower viscosities, characterized by greater nonlinearity, the M2NO-LPF variant proves more effective, especially at finer resolutions, underscoring its superior handling of more challenging dynamic conditions. This illustrates the robustness of both M2NO variants in adapting to diverse computational demands and physical scenarios.
D.3 2D Darcy Flow
Darcy flow [12, 30] is a fundamental relationship in hydrogeology, describing the flow of a fluid through a porous medium. We experiment on the steady-state of the 2D Darcy flow equation on the unit box, where it takes the following form:
$\nabla\cdot(a(x)\nabla u(x))$ $=f(x),$ $x$ $\in(0,1)^{2},$ (75)
$u(x)$ $=0,$ $x$ $\in\partial(0,1)^{2},$
where $a(x)$ is the permeability field, $u(x)$ is the hydraulic head, and $f(x)$ is the source term.
Results. Table 7 delineates the performance benchmarks on the Darcy Flow equation across varying input resolutions ($s$). Notably, both M2NO and M2NO-LPF exhibit competitive $L_{2}$ loss metrics, particularly excelling at lower resolutions. While M2NO maintains strong performance consistently across the range, M2NO-LPF demonstrates exceptional results at $s=512$, with an $L_{2}$ loss of 9.64E-03, suggesting enhanced efficiency in capturing flow characteristics at higher resolutions. These findings indicate the effectiveness of the M2NO models in accurately modeling fluid flow through porous media, which is central to the simulation of Darcy Flow.
Table 7: Benchmarks on Darcy Flow equation at various input resolution s.
Model s=32 s=64 s=128 s=256 s=512
Unet 8.43E-02 6.18E-02 5.71E-02 6.10E-02 9.25E-02
[…]
M2NO 4.17E-02 4.75E-02 4.85E-02 5.83E-02 2.73E-02
[…]
D.4 2D Diffusion-Reaction Equation
We consider a two-dimensional diffusion-reaction type PDE that combines a diffusion process with a rapid evolution from a source term [30]. The equation is written as:
$\partial_{t}u=D_{u}\partial_{xx}u+D_{u}\partial_{yy}u+R_{u},\quad\partial_{t}v=D_{v}\partial_{xx}v+D_{v}\partial_{yy}v+R_{v},$ (76)
where $D_{u}$ and $D_{v}$ are the diffusion coefficients for the activator and inhibitor, respectively, and $R_{u}=R_{u}(u,v)$ and $R_{v}=R_{v}(u,v)$ are the reaction functions for the activator and inhibitor, respectively.
The reaction functions are defined by the FitzHugh-Nagumo equation [19], written as:
$R_{u}(u,v)$ $=u-u^{3}-k-v,$ (77)
$R_{v}(u,v)$ $=u-v,$
where $k=5\times 10^{-3}$, and the diffusion coefficients for the activator and inhibitor are $D_{u}=1\times 10^{-3}$ and $D_{v}=5\times 10^{-3}$, respectively. The initial condition is generated as standard normal random noise $u(0,x,y)\sim\mathcal{N}(0,1.0)$ for $x\in(-1,1)$ and $y\in(-1,1)$.
We include the 2D diffusion-reaction equation as an example because it serves as a challenging benchmark problem. It involves two variables of interest, the activator and inhibitor, which are non-linearly coupled. Additionally, it has real-world applicability, particularly in modeling biological pattern formation [31, 30].
Table 8: Benchmarks on Diffusion-Reaction equation at various input resolution s.
Model $s=32$ $s=64$ $s=128$
Unet 2.18E-01 1.64E-01 1.46E-01
FNO 2.00E-01 1.56E-01 1.64E-01
MGNO 2.03E-01 1.50E-01 1.12E-01
MWT Leg 2.20E-01 1.51E-01 1.15E-01
LSM 2.92E-01 2.02E-01 1.58E-01
M2NO 1.97E-01 1.42E-01 1.01E-01
[…]
Promotion 1.60% 5.34% 10.15%
Results. Table 8 displays the performance benchmarks on the 2D Diffusion-Reaction equation, a complex model that simulates biological pattern formation involving non-linearly coupled variables—activator and inhibitor. Across various input resolutions ($s$), the M2NO models, particularly M2NO and M2NO-LPF, demonstrate improved performance compared to traditional methods such as U-net, FNO, and others. Notably, the M2NO model achieves the lowest $L_{2}$ losses at higher resolutions ($s=128$), indicating a 10.15% improvement in accuracy, reflecting its capability to more effectively handle the intricate dynamics of this challenging benchmark. The M2NO-LPF variant also shows significant performance enhancements, particularly at medium resolutions ($s=64$), suggesting its adeptness in capturing essential features of the diffusion-reaction processes.
D.5 Navier-Stokes Equation
Navier-Stokes equations describe the motion of viscous fluid substances [22, 12, 15], and are used to model ocean currents, weather, and air flow. We experiment on the 2D Navier-Stokes equation for a viscous, incompressible fluid in vorticity form on the unit torus, which takes the following form:
$\frac{\partial w(x,t)}{\partial t}+u(x,t)\cdot\nabla w(x,t)-\nu\Delta w(x,t)$ $=f(x),$ $\quad x$ $\in(0,1)^{2},$ $\quad t$ $\in(0,T],$ (78)
$\nabla\cdot u(x,t)$ $=0,$ $\quad x$ $\in(0,1)^{2},$ $\quad t$ $\in[0,T],$
$w_{0}(x)$ $=w(x,t=0),$ $\quad x$ $\in(0,1)^{2},$
where $w(x,t)$ is the vorticity, $u(x,t)$ is the velocity field, $\nu$ is the kinematic viscosity, and $f(x)$ is a forcing function.
Table 9: Benchmarks on Navier-Stoke equation at various $\nu$, $T$ and $N$.
Model $\nu$=1e-3 $\nu$=1e-4 $\nu$=1e-5
T=50 T=30 T=20
N=1000 N=10000 N=1000
UNet 3.86E-03 1.38E-02 2.15E-02
FNO-2D 4.95E-03 9.56E-03 1.72E-02
MGNO 3.39E-03 6.58E-03 1.27E-02
MWT-2D-Leg 1.83E-03 3.83E-03 1.19E-02
LSM 6.00E-03 1.02E-02 1.71E-02
[…]
Promotion 49.75% 58.47% 65.77%
Results. Table 9 displays the performance of various models on the Navier-Stokes equation across different viscosities, temporal resolutions, and grid sizes. The Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO) consistently outperforms other models such as UNet, FNO-2D, and LSM. M2NO achieves its lowest $L_{2}$ error of $9.17\times 10^{-4}$ under conditions of $\nu=1\times 10^{-3}$, $T=50$, and $N=1000$. This superior accuracy is maintained even under more challenging conditions; for instance, at $\nu=1\times 10^{-4}$ with $T=30$ and $N=10000$, it registers an $L_{2}$ error of $1.59\times 10^{-3}$, and at the finest tested viscosity of $\nu=1\times 10^{-5}$, $T=20$, and $N=1000$, the error is $4.07\times 10^{-3}$. M2NO not only shows remarkable stability and performance efficiency across these diverse settings but also demonstrates significant improvements in model accuracy, with promotion rates of 49.75%, 58.47%, and 65.77% across the various test conditions.
Appendix E Ablation and Hyperparameter Study
The ablation study uses Burgers’ equation with $\nu=1e-3$ under the resolution of 1024 as a reference example. The baseline M2NO is configured with levels $[1,1,1,2,2,2,3]$, $k=8$, $c=8$, Schur-based operators and comprises 5 architectural layers. We conduct the ablation and hyperparameters study by changing one model configuration and fixing the others.
Table 10: Ablation results comparing the $L_{2}$ error of different M2NO configurations. The removal of multilevel and multiwavelet features, along with a variant using a low-pass filter (M2NO-LPF), are evaluated against the baseline model.
Model Configuration $L_{2}$ Error
w/o multilevel 4.53E-05
w/o multiwavelet 4.56E-05
M2NO-LPF 7.52E-06
M2NO (Baseline) 7.16E-06
Ablation Study Analysis. In our ablation study of the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO) on Burgers’ equation with $\nu=1e-3$ and a resolution of 1024, summarized in Table 10, we investigated the impact of different model configurations on $L_{2}$ error. The baseline M2NO, utilizing full multilevel and multiwavelet features, achieved an $L_{2}$ error of 7.16E-06. Removal of the multilevel component or the multiwavelet feature resulted in similar increases in error to 4.53E-05 and 4.56E-05, respectively, underscoring their critical roles in reducing error. The M2NO-LPF variant, employing a low-pass filter instead of a traditional multigrid operator, demonstrated a slight increase in error to 7.52E-06, suggesting that while LPF simplifies the model, it retains much of the baseline’s efficacy, particularly in smoothing high-frequency details. These findings highlight the essential contributions of multilevel and multiwavelet components to the M2NO framework and suggest the potential for simpler models in specific scenarios.
Table 11: Hyperparameter study results for the M2NO model, detailing the $L_{2}$ error across various configurations. Changes in grid level, wavelet number, basis number, and layer number are evaluated against the baseline to assess their impact on model performance.
Hyperparameter Type Model Configuration $L_{2}$ Error
Iteration [1,1,1,1,1,1,1] 7.35E-06
Grid Level [1,2,3] 6.45E-06
Wavelet Number c=2 1.83E-05
c=4 8.16E-06
c=6 6.20E-06
c=16 5.29E-06
Basis Number k=2 1.28E-05
k=4 7.15E-06
Layer Number layer=3 8.24E-06
layer=4 7.19E-06
layer=6 4.66E-06
Baseline M2NO 7.16E-06
Hyperparameter Study Analysis. The hyperparameter study of M2NO, presented in Table 11, explored the impact of various model configurations on the $L_{2}$ error. Adjustments were made across multiple parameters, including iteration steps, grid levels, wavelet numbers, basis numbers, and the number of layers. Notably, increasing the grid levels from a flat [1,1,1,1,1,1] to a varied [1,2,3] configuration reduced the $L_{2}$ error from 7.35E-06 to 6.45E-06. Changes in the wavelet number revealed that using c=16 minimized the error to 5.29E-06, indicating a more significant wavelet count enhances accuracy. Similarly, an optimal layer count was identified, with six layers achieving the lowest error at 4.66E-06. Conversely, varying the basis number showed that a higher count does not always correlate with lower error, as k=4 performed better than k=2. These results illustrate the nuanced impacts of each parameter on model performance, guiding further refinement and optimization of M2NO configurations for enhanced precision in scientific simulations.
Appendix F Implementation Details
F.1 Baseline and Benchmark Construction
Time-dependent tasks. In our benchmarks, all tasks are time-dependent, except for Darcy Flow. These tasks are auto-regressive, requiring sequential predictions for each subsequent time step. Accordingly, we employ the 2D versions of the Multiwavelet-based Algebraic Multigrid Neural Operator (M2NO), along with the Fourier Neural Operator (FNO) and Multiwavelet Transform (MWT) models, to predict the next time frame in a step-by-step manner across all benchmarks.
Data splits. For the Diffusion-Reaction and Navier-Stokes (NS) datasets, the data is allocated with 80% used for training, 10% for validation, and 10% for testing. For all other datasets, the split is 60% for training, 20% for validation, and 20% for testing.
Baselines. We implemented all baseline models using their official source code and corresponding default parameters. However, for the LSM model, the default patch size of $[4,4]$ was not suitable for a resolution of $32\times 32$. Consequently, we adjusted the patch size to $[2,2]$ to ensure compatibility. For fairness in comparison, all methods were trained over 500 epochs using the $L_{2}$ loss function [22] and the ADAM optimizer [18] with an initial learning rate of $10^{-3}$, which was halved every 100 epochs ($\gamma=0.5$). Batch sizes were set at 512 for 1D PDEs and 16 for 2D PDEs. All experiments were conducted on a single Nvidia A800 80GB GPU.
F.2 Model Architecture
In the Multiwavelet-based Multigrid Neural Operator (M2NO), the neural networks $A$, $S$, $B$, $C$ and $D^{-1}$ are implemented as single-layer convolutional neural networks (CNNs). The orthonormal polynomial (OP) basis utilized is the Legendre (Leg) basis, chosen with both uniform and non-uniform measures $\mu_{0}$, in alignment with the Multiwavelet Transform (MWT)[12] methodologies.
From a computational perspective, the $restriction$ and $prolongation$ operators depicted in (16) and Fig 1 specifically target the transformation of multiscale and multiwavelet coefficients. Despite this focus on coefficients, the actual input and output of the model are point-wise function samples, specifically $(a_{i},u_{i})$. To address this, we synthesize the data sequence into conceptual functions $f_{a}=\sum_{i=1}^{N}a_{i}\phi_{ji}^{n}$ and $f_{u}=\sum_{i=1}^{N}u_{i}\phi_{ji}^{n}$, positing them within the space $V_{n}^{k}$ where $n=\log_{2}N$. It is critical to note that these functions, $f_{a}$ and $f_{u}$, are employed primarily for theoretical articulation and not for direct computation, serving as a notational convention rather than functional entities in the model’s operations.
Generalizing Dimensionality According to Discrete Wavelet Transform (DWT) theory, the dimension of the analyzed matrix must be a power of 2 [27]. Within the Algebraic Multigrid (AMG) context, such a restriction could severely limit the method’s applicability. Consequently, we generalize as follows: the dimension of the restriction operator $I_{h}^{2h}$, for all $k$, is defined to be $(N/2)\times N$ where
$N=\begin{cases}n,&\text{if }n\text{ is even},\\ n+1,&\text{if }n\text{ is odd},\end{cases}$ (79)
with $n=\text{dim}(A_{j})$. If $n$ is odd, the dimensions of the matrix and the corresponding right-hand side vector are incremented by one.
F.3 Model Configurations
The primary configurations for the Multiwavelet-based Multigrid Neural Operator (M2NO) are detailed in Table 12, and those for the M2NO with Low-Pass Filter (M2NO-LPF) in Table 13. Except where specifically noted, model parameters remain consistent across different hyperparameters and resolutions within the same benchmark.
Table 12: Model Configurations of M2NO.
Modules Advection Burgers Navier-Stokes Darcy Diffusion-Reaction
num_layer 5 5 5 5 5
c 8 8 2 1 4
k 8 8 8 4 4
grid_levels [1,1,2,2,2,3,3] [1,1,1,2,2,2,3] [1,1,2,2,2,3] [3,3,2,2,1] [2,2,1,1,1]
padding_mode ’circular’ ’circular’ ’zeros’ ’zeros’ ’zeros’
bias TRUE TRUE TRUE TRUE TRUE
base ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’
initializer ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’
norm TRUE TRUE TRUE TRUE TRUE
operator ’schur’ ’schur’ ’schur’ ’schur’ ’schur’
activation ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’
Table 13: Model Configurations of M2NO-LPF.
Modules Advection Burgers Navier-Stokes-1e-3 Navier-Stokes-1e-4 Navier-Stokes-1e-5 Darcy Diffusion-Reaction
num_layer 5 5 5 5 5 5 4
c 8 8 1 1 2 2 1
k 8 8 8 8 8 4 8
grid_levels [1,1,1,2,2,2,3] [1,1,1,2,2,2,3] [1,1,2,2,2,3] [1,1,2,2,2,3] [1,1,2,2,2,3] [3,3,2,2,1] [3,3,2,2,2]
padding_mode ’circular’ ’circular’ ’circular’ ’circular’ ’circular’ ’circular’ ’circular’
bias TRUE TRUE TRUE TRUE TRUE TRUE TRUE
base ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’ ’legendre’
initializer ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’ ’xavier_normal’
norm TRUE TRUE TRUE TRUE TRUE TRUE TRUE
operator ’lpf’ ’lpf’ ’lpf’ ’lpf’ ’lpf’ ’lpf’ ’lpf’
activation ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’ ’gelu’